OpenStudy (anonymous):

find all solutions to the equation 2cos^2x=13sinx-5 in the interval (0,2pi)

OpenStudy (anonymous):

First step is to replace cos^2x with 1 - sin^2x

OpenStudy (anonymous):

then waht?

OpenStudy (anonymous):

2(1 - sin^2x) = 13sinx - 5 2 - 2sin^2x = 13sinx - 5 0 = 2sin^2x + 13sinx - 7 Now factor it like a quadratic 0 = (2sinx - 1)(sinx + 7) 2sinx - 1 = 0 or sinx + 7 = 0. sinx + 7 can't = 0 since sinx can't = -7.. So that leaves us with 2sinx - 1 = 0 2sinx = 1 sinx = 1/2 x = pi/6 and 5pi/6

OpenStudy (anonymous):

oh i see thanks